An investment will pay you $22,000 in 7 years. The appropriate discount rate is 10 percent compounded daily.
Required:
What is the present value?


rev: 09_17_2012

Multiple Choice

  • $11,253.70

  • $11,472.22

  • $10,925.92

  • $10,379.63

  • $11,289.48
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

22000=P*(1+0.1/365)^(365*7)

P=22000/(1+0.1/365)^(365*7)

=22000*0.496632915

=$10925.92(Approx).